Mathematics 243 – Algebraic Structures Definitions Quiz 6 – October 26, 2012

Your Name:

A) (5) If a ≡ 8341 (mod 100) and b ≡ 1702 (mod 100), find r in the range 0 ≤ r < 100 such that ab ≡ r (mod 100).

B) (5) Define: f : A → B is onto.